OpenStudy (anonymous):

Translate the phrase into a math expression. 12 less than twice a number A. 2 · (n – 12) B. 2 · n - 12 C. 2 · n ÷ 12 D. 12 – 2 · n

OpenStudy (sepeario):

They want the number here to be n. Twice this is 2*n=2n, and subtract makes it 2n-12. Hence B.

OpenStudy (unklerhaukus):

\[ \color{brown}{\overbrace{\text{12 less than }}^{-12}} \color{green}{\overbrace{\text{twice }}^{2\cdot}} \color{teal }{\overbrace{\text{a number}}^n}\\ =\color{brown}{-12} +\color{green}{2\cdot} \color{teal} n\\\hspace{6em} \textit{<rearranging>}\\ =\color{green}{2\cdot}\color{teal} n\color{brown}{-12}\]

OpenStudy (skullpatrol):